10. Problem 4.10 (M-B, SHARE PRICE, and EV-EBITDA) 

You are given the following information: Stockholders' equity as reported on the firm's balance sheet = $3.25 billion, price/earnings ratio = 22, common shares outstanding = 180 million, and market/book ratio = 1.6. The firm's market value of total debt is $6 billion, the firm has cash and equivalents totaling $250 million, and the firm's EBITDA equals $1 billion. What is the price of a share of the company's common stock? Do not round intermediate calculations. Round your answer to the nearest cent. 

What is the firm's EV/EBITDA? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Enterprise Value = Market capital for equity + Value of debt + Cash & cash equivalent

Market value / book ratio  = 1.60

Market value = 1.60 * book value = 1.6 * 3.25 = 5.2 billion = 5200 million

Price per share = Market value of equity / no of shares outstanding

= 5200 / 180 =28.89

Enterprise Value = 5200 + 6000 + 250 = 11450 million

EV / EBITDA = 11450 / 1000 = 11.45